OpenStudy (anonymous):

f(x)=sqrt{ln(3x ^{4}}-2x) help plzz

OpenStudy (anonymous):

\[f(x)=\sqrt{\ln(3x ^{4}}-2x)\]

OpenStudy (rsadhvika):

what you need help with ?

OpenStudy (anonymous):

derived

OpenStudy (anonymous):

well u need to apply chain rule here when u have a composite function like this \( f(x)=g(h(x)) \) then u can write \( f'(x)= h'(x) \ g'(h(x)) \) for our particular case \( g(x)=\sqrt{x} \) and \( h(x)=\ln (3x^4-2x) \) and use this properties \(\large \frac{d}{dx} \sqrt{u(x)}=\frac{u'(x)}{2\sqrt{u(x)}} \) and \(\large \frac{d}{dx} \ln [u(x)]=\frac{u'(x)}{u(x)} \)

OpenStudy (anonymous):

i don't understand:(

OpenStudy (anonymous):

my explanation is not well :(

OpenStudy (anonymous):

yeah..its difucult

OpenStudy (anonymous):

@sami-21

OpenStudy (phi):

if you take the derivative of \[ f= \sqrt{u -2x} \] you get \[ \frac{df}{dx}= \frac{1}{2}u^{-\frac{1}{2}} \frac{du}{dx} \] if u happens to be \[ u= ln(3x^4-2x) \] can you find du/dx ?

OpenStudy (phi):

* make the first equation \[ f= \sqrt{u} \] not sqrt(u-2x)

OpenStudy (anonymous):

Oh Oh... o.O \( f(x)=\sqrt{\ln (3x^4)}-2x \) ?

OpenStudy (phi):

not sure, but I am guessing it's ln(3x^4-2x) all under the square root.

OpenStudy (anonymous):

Applying the derivatives separately if question given by Kazem is right then..

OpenStudy (anonymous):

But I think that whole will be under the root..

OpenStudy (usukidoll):

ugh I hate my pc. I was drawing it, but it got messed up.

OpenStudy (anonymous):

oo

OpenStudy (usukidoll):